find the 31st term of the aritmctic sequence 23,36,49,.....
Answer by stanbon(75887)

find the 31st term of the aritmctic sequence 23,36,49,.....
---
a(1) = 23
d = 36-23 = 13
----------------------
Form: a(n) = 23+(n-1)d
---
a(31) = 23 + 30*13
a(31) = 23 + 390
a(31) = 413
==========================
Cheers,
Stan H.
==========================